# Couples can be tractable: New algorithms and hardness results for the Hospitals / Residents problem with Couples

### v2

@@ 4.3 Matchings with the minimum number of blocking pairs
We have established that if $I$ admits a complete stable matching then $J$ admits a matching with at most $n_{t}$ blocking pairs. An $m_{I}^{C}$-approximation algorithm for min bp hrc will thus return a matching in $J$ with at most $n_{t}m_{I}^{C}$ blocking pairs. On the other hand if $I$ admits no complete stable matching then any matching in $J$ has at least $B=n_{t}m_{I}^{C}+1$ blocking pairs, so any approximation algorithm for min bp hrc will return a matching in $J$ with at least $n_{t}m_{I}^{C}+1$ blocking pairs.
@@
Hence the existence of an $m_{I}^{C}$-approximation algorithm for min bp hrc implies a polynomial-time algorithm for com smti in a special case that is NP-complete, a contradiction unless P=NP. Let $m$ denote the total length of the hospitals’ preference lists in $J$. We claim that $m_{I}^{C}\geq m^{1-\varepsilon}$, and then the proof is complete.
@@
$m_{I}^{C}$$\geq$ $7^{-\frac{C}{C+2}}m^{\frac{C}{C+2}}\geq m^{\frac{C-1}{C+2}}=m^{1-\frac{3}{C+2}}=m^{1-\varepsilon}$ (3)
@@ 5 Conclusions and open problems
In this paper, we have presented a range of novel polynomial-time solvability and NP-hardness results for variants of hrc, with a heavy focus on the natural restriction that the couples’ preferences are sub-responsive and sub-complete. The results in this paper contribute to building knowledge about the frontier between polynomial-time solvable and NP-hard variants.

### v3

@@ 4.3 Matchings with the minimum number of blocking pairs
We have established that if $I$ admits a complete stable matching then $J$ admits a matching with at most $n_{t}$ blocking pairs. An $m_{I}^{C}$-approximation algorithm for min bp hrc will thus returns a matching in $J$ with at most $n_{t}m_{I}^{C}$ blocking pairs. On the other hand if $I$ admits no complete stable matching then any matching in $J$ has at least $B=n_{t}m_{I}^{C}+1$ blocking pairs, so any approximation algorithm for min bp hrc will return a matching in $J$ with at least $n_{t}m_{I}^{C}+1$ blocking pairs.
@@
Hence, the existence of an $m_{I}^{C}$-approximation algorithm for min bp hrc implies a polynomial-time algorithm for com smti in a special case that is NP-complete, a contradiction unless P=NP. Let $m$ denote the total length of the hospitals’ preference lists in $J$. We claim that $m_{I}^{C}\geq m^{1-\varepsilon}$, and then the proof is complete.
@@
$m_{I}^{C}$$\geq$ $7^{-\frac{C}{C+2}}m^{\frac{C}{C+2}}\geq m^{\frac{C-1}{C+2}}=m^{1-\frac{3}{C+2}}\geq m^{1-\varepsilon}$ (3)
@@ 5 Conclusions and open problems
In this paper, we have presented a range of novel polynomial-time solvability and NP-hardness results for variants of hrc, with a heavy focus on the restriction that the couples’ preferences are sub-responsive and sub-complete. The results in this paper contribute to building knowledge about the frontier between polynomial-time solvable and NP-hard variants.
@@ Acknowledgements
We would like to thank the anonymous reviewers of the preliminary version of this paper submitted to IJCAI 2024 for their valuable comments, which helped to improve the presentation of this paper.
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.